OpenStudy (anonymous):

need super help find the exact value of cos theta/2 given... sintheta= 5/13,90degrees

OpenStudy (anonymous):

c since sin theta = 5/13, then cos theta = 12/13; Use the cos theta/2 identity to solve this

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

cos theta/2 = +- sqrt ((1 + cos theta)/2) = sqrt ((1 + 12/13)/2) sqrt (25/13 /2) = sqrt (25/26) = 5/sqrt 26 = 6 root 26/26